Suppose the supply of lemonade is given by Q= 40p were q is measured in pints and p is measured in cents per pint and the demand for lemonade is given by Q=5000-10p. Now suppose bumper weather conditions improve the harvest of lemons so producers increase the quantity supplied of lemonade by 200 pints at every price. What are the equilibrium P and Q for a lemonade after the bumper weather conditions


1
Expert's answer
2022-02-20T17:05:30-0500

"Quantity demanded=Quantity supplied"

"5000-10p=40p"

"5000=50p"

"p=100cents"

"Q=40*100"

"Q=4000"pints before increase in "QS"

since the quantity supplied has increased by 20 hence, "QS" is;

"Q=40*200P" but P is 100 cents

"Q=40*200*100"

"Q=800000 pints"